# Understanding Long Videos in One Multimodal Language Model Pass

### v1

@@ Abstract
Large Language Models (LLMs), known to contain a strong awareness of world knowledge, have allowed recent approaches to achieve excellent performance on Long-Video Understanding benchmarks, but at high inference costs. In this work, we first propose Likelihood Selection, a simple technique that unlocks faster inference in autoregressive LLMs for multiple-choice tasks common in long-video benchmarks. In addition to faster inference, we discover the resulting models to yield surprisingly good accuracy on long-video tasks, even with no video specific information. Building on this, we inject video-specific object-centric information extracted from off-the-shelf pre-trained models and utilize natural language as a medium for information fusion. Our resulting Multimodal Video Understanding (MVU) framework demonstrates state-of-the-art performance across long-video and fine-grained action recognition benchmarks. Code available at: https://github.com/kahnchana/mvu
Keywords:
Video Understanding Multi-Modal Language Models
@@ 1 Introduction
Figure 1: Overview of Framework: We propose three variants of our framework that solves complex long-video question-answering tasks. (left-top) Just LLM utilizes only world knowledge with zero task-specific awareness. (left-right) Single Frame VLM processes an additional center frame to obtain task context but accesses no video specific information. (right) Our complete approach, MVU extracts three additional object-centric modalities followed by fusion in language space. LS refers to likelihood selection.
@@
Videos contain unique forms of information beyond what is contained within a single natural image [7]. The recent success of large language models (LLMs) and their visual extensions, Vision-Language Models (VLMs), have resulted in incredible performance on a range of complex language-tied video understanding benchmarks [83, 35, 77, 3], particularly on long-video question answering: a task that requires awareness over longer temporal windows [35] as well as causal and temporal action reasoning [65]. However, the LLMs underlying these approaches contain extensive world knowledge (e.g. understanding of physics, culture, human common-sense [76, 1, 85, 63, 70, 30]), raising the question of whether they excel at video tasks due to actual video modality awareness or simply utilizing world knowledge and contextual information. Such understanding of model reasoning is important for robust deployments avoiding spurious correlation based predictions [80, 49] as well as for model interpretability [51, 25].
@@
We systematically study this question in the context of long-video question-answering benchmarks. First, motivated by the next-token prediction objective in LLM training [45] and the choice selection aspect of long-video benchmarks, we propose an efficient inference strategy utilizing intermediate logits of autoregressive models extracted from a single forward pass, termed likelihood selection. By doing so, we unlock a faster inference ability for autoregressive LLMs/VLMs, that can generate selection results in one forward pass. We then construct two modality-constrained frameworks tagged Just LLM and Single Frame VLM. The first is constrained to access only world knowledge (i.e. no task-specific information beyond the question), whereas the latter has access to task context with an additional single center-frame input. We discover how these models perform significantly better than a random prediction on long-video understanding benchmarks (see Tab. 4). In fact, the latter utilizing purely world knowledge and contextual information outperforms multiple recent state-of-the-art video understanding works (see Tab. 1), challenging the notion of how much video is actually utilized by existing approaches to solve these complex video tasks.
@@
Motivated by these findings, we study how to inject additional video-specific information into our preliminary frameworks using natural language in an interpretable fashion, and whether it can improve their video understanding. We explore three forms of object-centric information modalities, develop pipelines requiring zero video-level training to extract such information, and utilize natural language to fuse this multi-modal information. Our resulting framework, tagged MVU, while achieving state-of-the-art zero-shot performance across long-video understanding benchmarks, also exhibits better interpretability (e.g. exposing video-specific information utilized) through its language-based operation.
@@
    Likelihood Selection: LLM-based choice selection in a single forward pass
@@
    Uncovering surprisingly strong performance on complex video-language tasks by resulting modality-constrained baselines
@@
    MVU framework for language-based fusion of video-specific modalities.
@@ 2 Related Work
Video Modality Exploration: Multiple recent works dissect video modality into individual components [80, 7, 50, 48]. Single frame baselines are one interesting sub-class [7, 13, 84, 54, 5]. Extracting object-centric video modalities is another idea, spanning back to [13] which extracts multiple small objects from frames followed by modeling relations across frames and objects. Similarly, [54, 84] combines spatial information with single images to perform video tasks. However, these prior approaches focus on simple video tasks (i.e. action recognition) limited to visual modality. In contrast, our approach tackles the more complex language-tied task of long-video question answering. This task is also explored in [7], but we differ with likelihood selection and multi-modal information fusion.
@@
Long Video Question Answering: Long-video question-answering benchmarks are constructed to specifically test strong causal and temporal reasoning [65] over long temporal windows [35]. Early works explore querying objects or events based on referential and spatial relations [69, 82, 79] followed by focus on temporal modeling of sequential events [27, 28, 20, 66, 67]. More recent works leverage large language models (LLMs) [77, 43, 62, 3, 73, 75] to directly perform these tasks but require video-caption training. In contrast, our MVU operates zero-shot on these tasks requiring no video-level training. Limited zero-shot operation is explored in [61, 83], but we differ in using object-centric modality information and efficient LLM sampling.
@@
Large Language Model Reasoning: Recent LLMs [41, 10, 9] demonstrate multiple forms of strong reasoning abilities [26, 11, 33] including combining different information [64]. Their recent open-source variants [59, 58, 21] achieve equally promising skills using scaled-down models [21] while also demonstrating strong world knowledge [76, 1, 85, 63, 70, 30]. In our work, we leverage these strengths of LLMs for complex video-language tasks, focused on disentangling the effect of their world knowledge from task context and video-specific information.
@@
Language based Fusion: The idea of fusing different modality information using natural language as a medium has been explored in multiple recent works [51, 31, 17, 60, 18, 81]. In [51, 31] language is utilized as an implicit medium for self-supervising video action recognition. Multimodal information represented as language is fused with visual information for action recognition tasks in [17, 60, 18]. We utilize a similar language-as-a-medium fusion of multimodal information but explore this in the context of complex video-language tasks. Work in [81] exploring image and video question-answering tasks is most similar to our work. Our proposed MVU, however, differs with a focus on long-video understanding, object-centric information modalities, and our LLM-specific proposals for efficient selection.
3 Solve in One LLM Pass
In this section, we first present Likelihood Selection, an efficient technique that unlocks autoregressive LLMs to solve multiple selection problems with only one independent forward pass. This is in contrast to next token sampling requiring iterative generations dependent on each previous output. Based on that, in Sec. 4, we introduce Multimodal Video Understanding (MVU), a simple framework that aggregates multimodal video-relevant information using natural language, which achieves significant improvements over baselines across multiple datasets.
@@ 3.1 Problem Formulation
We focus on two types of video understanding tasks:
@@
    Fine-Grained Video Action Recognition (N-Way Classification)
and construct a unified problem formulation for both tasks aforementioned.
Consider a video $x_{v}\in\mathbb{R}^{L\times H\times W\times C}$, a textual question $x_{t}$, a set of textual candidate answers $Y=\left\{y_{i},i=1,...,M\right\}$, and a model $V(\cdot)$ selecting one answer from the given set of answers (noted as $\hat{y}:=V(x_{v},x_{t},Y)$). Selected $\hat{y}$ should ideally be identical to groundtruth $y_{g}$. Here $L,H,W,C$ are the number of frames of the video, frame height, width, and number of channels respectively. $M$ is the number of candidate answers. For multiple-choice-based selection tasks, $x_{v}$, $x_{t}$, and $Y$ are directly present in dataset. For N-Way Classification tasks, we set $x_{t}$ as a generic question (details in Appendix 0.A) and formulate $Y$ by applying a fixed template to the labels of all N classes of the dataset. This formulation is used for the remainder of the paper unless a specific exception is noted.
@@ 3.2 Likelihood Selection
Recent autoregressive LLMs and their visual extensions (V-LLMs or VLMs) possess extensive world knowledge that is highly generic and transferable. They are commonly used through iterative prediction of the next tokens conditioned on prior outputs to generate complete natural language outputs. Such a generation process is usually modeled as sampling from a conditional likelihood shown as  Eq. 1, where $\hat{y}^{j}$ stands for the $j^{\mathrm{th}}$ token in a textual sequence $\hat{y}$ autoregressively generated by the model.
$P(\hat{y}|x_{t})$ $=\prod_{j}P(\hat{y}^{j}|\hat{y}^{1,...,j-1},x_{t})$ (1)
$P(y_{i}|x_{t})$ $=\prod_{j}P(y^{j}_{i}|y^{1,...,j-1}_{i},x_{t})$ (2)
Assume a perfect LLM, intuitively when $y_{i}$ is a proper answer to the question $x_{t}$ (say $y_{i}=y_{g}$), the conditional likelihood $P(y_{i}|x_{t})$ should be larger than any other $P(y_{w}|x_{t})$ where $y_{w}$ is a wrong answer to question $x_{t}$. In fact, modern LLMs are trained with a similar objective [45]. Motivated by this standard LLM training objective, we estimate the relative numerical scales of conditional likelihood on different answers $P(y_{i}|x_{t})$ using a cross-entropy error $e_{i}$, given their equivalence (negative log-likelihood and multiclass cross-entropy, see Section 4.3.4 in [6]). We calculate $e_{i}$ with a single forward pass of LLM without detokenization and the selection can be made by simply picking up the answer with the lowest error, equivalent to the highest conditional likelihood among all the answers.
@@
Figure 2: Likelihood Selection Workflow: We illustrate how our proposed selection strategy can be efficiently parallelized and calculated with a simple cross-entropy loss in one forward pass, followed by an argmin operation, without generating additional tokens. See Sec. 3 for more details.
Therein, we propose Likelihood Selection, first illustrated with a toy example in Fig. 2, where the task is vanilla question-answering with only textual context and the model takes one question $x_{t}$ as well as $M=5$ candidate answers $y_{1,...,5}$. To find the best answer, we simply concatenate the question with each candidate independently ($s_{i}=\mathrm{concat}\left(x_{t},y_{i}\right)$ ) and pad them into a batch $\{s_{1,...,5}\}$. Then the LLM takes the batch of five samples with casual attention masks and performs one inference forward pass, resulting in five shifted logits $\{p_{1,...,5}\}$. Next, we shift the input sequence $s_{i}$ to align the logits $p_{i}$ and calculate the average cross-entropy error only on tokens of $y_{i}$ Finally, the answer with the smallest $e_{i}$ will be picked up as the best answer. The method can be formulated as in Eq. 4 using equivalence of negative log-likelihood to cross-entropy in Eq. 3. Here $n_{i}$ stands for the token sequence length of $y_{i}$ and $p_{i}^{j}$ stands for logits of the $j^{\mathrm{th}}$ token in $p_{i}=V(\mathrm{concat}\left(x_{t},y_{i}\right))$ with logits limited to only those of $y_{i}$.
In summary, Likelihood Selection performs one single forward pass to extract the network logit outputs, calculates error ($e_{i}$) on each choice, and selects the choice with the lowest error. Note that our method does not utilize any iterative autoregressive generation using the LLM. This results in considerable speed-ups for inference time. We also obtain the additional advantages of avoiding LLM hallucinations and deviations from expected output formats over iterative generation strategies applied to similar visual tasks [18] leading to better accuracy (see Tab. 6.). In Sec. 5.3, we demonstrate both our speed gains and performance improvements.
Furthermore, Likelihood Selection is a generic method that can be easily extended to autoregressive VLMs, and in principle, there is no reason it could not also be used with extra modalities besides language. In the coming section, we will extend this method to a VLM and unlock its ability to process token sequences from multimodalities in one forward pass.

### v2

# Understanding Long Videos with Multimodal Language Models

@@ Abstract
Large Language Models (LLMs) have allowed recent LLM-based approaches to achieve excellent performance on long-video understanding benchmarks. We investigate how extensive world knowledge and strong reasoning skills of underlying LLMs influence this strong performance. Surprisingly, we discover that LLM-based approaches can yield surprisingly good accuracy on long-video tasks with limited video information, sometimes even with no video specific information. Building on this, we exploring injecting video-specific information into an LLM-based framework. We utilize off-the-shelf vision tools to extract three object-centric information modalities from videos and then leverage natural language as a medium for fusing this information. Our resulting Multimodal Video Understanding (MVU) framework demonstrates state-of-the-art performance across multiple video understanding benchmarks. Strong performance also on robotics domain tasks establish its strong generality. Our code will be released publicly.
@@ 1 Introduction
Recent success of large language models (LLMs) and their visual extensions, vision-language models (VLMs), has led to incredible performance on complex language-tied video understanding benchmarks (Zhang et al., 2023a), particularly on long-video question answering: a task that requires awareness over longer temporal windows (Mangalam et al., 2023) as well as causal and temporal action reasoning (Xiao et al., 2021). However, the LLMs underlying these approaches contain extensive world knowledge (e.g. understanding of physics, culture, human common-sense) and reasoning abilities (Yu et al., 2023a; Wang & Zhao, 2023), raising the question of whether they excel at video tasks due to actual video modality awareness or simply utilizing world knowledge and contextual information. Such understanding of model reasoning is important for robust deployments avoiding spurious correlation based predictions as well as for better model interpretability (Yun et al., 2022; Xiao et al., 2024).
@@
In this work, we systematically study this question in the context of video question-answering (QnA) benchmarks, building two modality-constrained baselines to highlight our findings. These two frameworks are tagged Just-LLM and Single-Frame-VLM. The first is constrained to access only the task textual query allowing it to answer using only world knowledge (i.e. no task-specific visual information). The latter is given access to task context with an additional single center-frame from the video as input. We discover how these models perform significantly better than random prediction on multiple long-video understanding benchmarks (see Table 1). In fact, the latter utilizing purely world knowledge and contextual information even outperforms multiple recent state-of-the-art video understanding works (see Table 2), challenging the notion of how much video information is actually utilized by existing approaches to solve these complex video QnA tasks.
We next focus on efficient inference to allow rapid experimentation with our LLM based frameworks. Therein, we explore suitable prompting and templating to adapt likelihood selection techniques from prior work (Robinson et al., 2023) to video QnA tasks. Our resulting framework achieves more efficient inference with competing performance in comparison to prior work that commonly use auto-regressive generation to tackle long-video QnA benchmarks (Zhang et al., 2023a; Balavzevi’c et al., 2024).
@@
Motivated by our initial findings on modality-constrained performance, we study how to inject additional video-specific information into our framework using natural language in an interpretable manner in order to further improve video understanding. We explore three forms of object-centric information modalities, develop pipelines requiring zero video-level training to extract such information using off-the-shelf vision tools, and utilize natural language to fuse this multi-modal information using templating operations. Our resulting approach, termed Multi-Modal Video Understanding (MVU) framework, while achieving state-of-the-art zero-shot performance across long-video understanding benchmarks, also exhibits better interpretability (e.g. exposing video-specific information utilized) through its language-based operation. Moreover, our proposed MVU exhibits generality with its strong performance even on robotics domain tasks.
@@
    Uncover surprisingly strong performance on complex video-language tasks by modality-constrained baselines with no access to video-specific information.
@@
    Adapting Likelihood Selection strategies to video QnA benchmarks for efficient evaluation.
@@
    Propose a novel VLM-based video QnA framework that extracts video specific object-centric information followed by natural language based fusion.
Extensive evaluation of our framework on video QnA benchmarks as well as robotics domain tasks establish its clear benefit. We also highlight how our proposed MVU is a framework (as opposed to a fixed model) that can easily be integrated with any newer, stronger LLMs or VLMs to utilize their improved capabilities.
Figure 1: Overview of Framework: We propose three variants of our framework that solves complex long-video question-answering tasks. (left-top) Just-LLM utilizes only world knowledge with zero task-specific awareness. (left-right) Single-Frame-VLM processes an additional center frame to obtain task context but accesses no video specific information. (right) Our complete approach, MVU extracts three additional object-centric modalities followed by fusion in language space. LS refers to likelihood selection.
@@ 2 Related Work
Video Modality Exploration: Multiple recent works dissect video modality into individual components (Yun et al., 2022; Buch et al., 2022; Ranasinghe et al., 2021; Ramasinghe et al., 2018). Single frame baselines are one interesting sub-class (Buch et al., 2022; Davis & Bobick, 1997; Zhao et al., 2017; Safaei & Foroosh, 2019; Bilen et al., 2016). Extracting object-centric video modalities is another idea, spanning back to Davis & Bobick (1997) which extracts multiple small objects from frames followed by modeling relations across frames and objects. Similarly, Safaei & Foroosh (2019); Zhao et al. (2017) combine spatial information with single images to perform video tasks. However, these prior approaches focus on simple video tasks (i.e. action recognition) limited to visual modality. In contrast, our approach tackles the more complex language-tied task of long-video question answering that necessitates strong causal and temporal reasoning over long temporal windows. This task is also explored in Buch et al. (2022), but we differ with likelihood selection, multi-modal information fusion, and usage of modern LLMs.
@@
Long Video Question Answering: Long-video question-answering benchmarks are constructed to specifically test strong causal and temporal reasoning (Xiao et al., 2021) over long temporal windows (Mangalam et al., 2023). Early works explore querying objects or events based on referential and spatial relations (Xu et al., 2017; Zeng et al., 2017; Yu et al., 2019) followed by focus on temporal modeling of sequential events (Lei et al., 2018; 2020; Hosseini et al., 2022; Xiao et al., 2022a; b). While motivated by these works, MVU integrates such object information with large language models (LLMs) in a zero-shot manner requiring no video-level training. More recent works leverage LLMs (Yu et al., 2023b; Papalampidi et al., 2023; Wang et al., 2022; Balavzevi’c et al., 2024; Yang et al., 2022; Ye et al., 2023b) to directly perform these tasks but require video-caption training. In contrast, our MVU operates zero-shot on these tasks requiring no video-level training. Limited zero-shot operation is explored in Wang et al. (2023); Zhang et al. (2023a), but we differ in using object-centric modality information, efficient LLM sampling, and zero video level training.
@@
Large Language Model Reasoning: Recent LLMs (OpenAI, 2023; Chowdhery et al., 2022; Chiang et al., 2023) demonstrate multiple forms of strong reasoning abilities (Kıcıman et al., 2023; Creswell & Shanahan, 2022; Liu et al., 2023b) including combining different information (Weston & Sukhbaatar, 2023). Their recent open-source variants (Touvron et al., 2023; Team et al., 2023; Jiang et al., 2023) achieve equally promising skills using scaled-down models (Jiang et al., 2023) while also demonstrating strong world knowledge (Yu et al., 2023a; AlKhamissi et al., 2024; Zhao et al., 2023; Wang & Zhao, 2023; Xu et al., 2024; Li et al., 2023d) even in domains such as robotics Li et al. (2024). In our work, we leverage these strengths of LLMs for complex video-language tasks, focused on disentangling the effect of their abilities for video QnA tasks.
@@
Language based Fusion: The idea of fusing different modality information using natural language as a medium has been explored in multiple recent works (Ranasinghe & Ryoo, 2023; Lin et al., 2023b; Hanu et al., 2022; Wang & Chen, 2017; Hanu et al., 2023; Zeng et al., 2022). In Ranasinghe & Ryoo (2023); Lin et al. (2023b) language is utilized as an implicit medium for self-supervising video action recognition. Multimodal information represented as language is fused with visual information for action recognition and robotics tasks in (Hanu et al., 2022; Wang & Chen, 2017; Hanu et al., 2023; Li et al., 2024). We utilize a similar language-as-a-medium fusion of multimodal information but explore this in the context of complex video-language tasks. Zeng et al. (2022) is most similar to our work, but we differ with focus on long-video tasks and object-centric information.
3 Naive Baselines & Likelihood Selection
In this section, we first establish our problem setting, then discuss adapting likelihood selection for video QnA tasks, and finally introduce two naive LLM based frameworks for video question answering tasks, tagged Just-LLM and Single-Frame-VLM (see Figure 1).
@@ 3.1 Problem Formulation
We focus on two categories of video understanding tasks:
@@
    Open Ended Video Question Answering (Text Generation)
For the first task, we construct a unified problem formulation accounting their choice based selection aspect. For the latter, we resort to standard LLM based answer generation.
Consider a video $x_{v}\in\mathbb{R}^{L\times H\times W\times C}$, a textual question $x_{t}$, a set of textual candidate answers $Y=\left\{y_{i},i=1,...,M\right\}$, and a model $V(\cdot)$ selecting one answer from the given set of answers (noted as $\hat{y}:=V(x_{v},x_{t},Y)$). Selected $\hat{y}$ should ideally be identical to groundtruth $y_{g}$. Here $L,H,W,C$ are the number of frames of the video, frame height, width, and number of channels respectively. $M$ is the number of candidate answers. For multiple-choice-based selection tasks, $x_{v}$, $x_{t}$, and $Y$ are directly present in dataset. For N-Way Classification tasks, we set $x_{t}$ as a generic question (details in Appendix A) and formulate $Y$ by applying a fixed template to the labels of all N classes of the dataset. This formulation is used for the remainder of the paper unless a specific exception is noted.
In the case of open ended video question answering, we follow standard settings of LLM based text generation for video tasks following Maaz et al. (2023).
@@ 3.2 Likelihood Selection
The common technique for LLM based approaches tackling question answering (QnA) tasks is likelihood based choice selection (also referred as Cloze Prompting, see Robinson et al. (2023)). Adopting such likelihood based selection for different tasks (or to VLMs) is however not straightforward (Robinson et al., 2023), leading to most existing long video QnA approaches resorting to LLM based answer generation. In fact, most existing long-video QnA approaches using LLMs / VLMs for choice selection (Papalampidi et al., 2023; Wang et al., 2022; Balavzevi’c et al., 2024) resort to full answer generation followed by embedding or template based matching to ground-truth choices, incurring significant inference costs for evaluation.
@@
Figure 2: Likelihood Selection Workflow: We illustrate how our proposed selection strategy can be efficiently parallelized and calculated with a simple cross-entropy loss in one forward pass, followed by an argmin operation, without iteratively generating multiple tokens.
In light of this, we explore prompting and likelihood calculation techniques optimal for applying Likelihood Selection on long video QnA tasks with either LLMs or even VLMs. Adapting this technique unlocks autoregressive LLMs / VLMs ability to solve multiple selection problems with only one forward pass. This is in contrast to next token sampling requiring iterative generations dependent on previous outputs for each answer token. As illustrated in Figure 2, likelihood selection combines each question $x_{t}$ with all $M=5$ candidate answers $y_{1,...,5}$ to calculate the conditional likelihood $P(y_{i}|x_{t})$ for each answer. This is calculated in a single LLM forward pass and selects the answer with the highest conditional likelihood among all the answers. We direct the reader to complete details of our video QnA specific prompting templates and calculations in Table A.4.
In addition to the considerable inference speed-up from likelihood selection, we also obtain the additional advantages of avoiding LLM hallucinations and deviations from expected output formats over iterative generation strategies applied to similar visual tasks (see Hanu et al. (2023)). We empirically validate the improved performance from such behaviour in our ablations (see Tab. 6.).
3.3 Modality Constrained Variants
We next introduce the two modality-constrained variants of our framework tagged Just-LLM and Single-Frame-VLM (illustrated in Figure 1). The former utilizes only the task question injected as language ($x_{t}$) with no other task-specific information. Note how this naive variant does not access any information extracted from the video for each task instance. The latter utilizes an additional center visual frame ($x_{v}^{c}$), extracted from the center of the video ($x_{v}$) timeline. This variant accesses no video-specific data (e.g. temporal or motion information). The center frame usage ensures no temporal information leakage in frame selection for this variant.
We hypothesize that Just-LLM with no access to task-specific knowledge is constrained to generate predictions utilizing its internal world knowledge (e.g. physics, culture, human common sense). We refer to this as world modality. For a given question regarding a natural video and a set of candidate answers, there is a possibility that one choice is more probable given how our world operates. In cases that this choice turns out to be correct, the internal world knowledge of the LLM allows it to easily select that choice resulting in above random performance. This variant of our framework highlights such cases in long video QnA tasks.
In the case of Single-Frame-VLM, it is provided with task information but is limited to a single frame, which could possibly provide important scene context. Therein, we refer to this variant as operating with world and contextual information modalities. For example, consider a video with a man walking a dog. The scene context of the dog and man combined with the LLM world knowledge and reasoning skills may be sufficient to correctly answer the question with no temporal or motion information. Performance of this variant highlights the prevalence of similar cases in long video QnA tasks when using LLM based approaches.
We evaluate these two modality-constrained variants and summarize our findings in Table 1. We uncover surprisingly strong performance of both variants on two long-video understanding benchmarks. In the case of Just-LLM variant, we achieve performance significantly higher than random selection (+25.8% on EgoSchema-S / +20.1% on NextQA-T) using zero visual information. This indicates the large portion of questions in existing video-QnA benchmarks that can be answered correctly purely using world knowledge. We also highlight our Single-Frame-VLM performing on par with state-of-the-art LLM based approach from Zhang et al. (2023a). In particular, for EgoSchema-S we outperform Zhang et al. (2023a) which uses information extracted from 180 frames per video incurring an inference cost over 100 times higher than ours. In light of these findings, we argue that long video understanding approaches in particular must focus on learning information beyond what a single frame baseline can achieve, possibly in an interpretable manner.
Table 1: Modality Constrained Variants: We evaluate on ES-S (public subset of EgoSchema (Mangalam et al., 2023)) and NextQA-T (test set of NextQA (Xiao et al., 2021)) datasets. We report accuracy (%) and inference time per sample (s). Recent state-of-the-art approach from Zhang et al. (2023a) using multi-frame inputs is also reported for comparison.
Method Param ES-S NextQA-T
Acc Time Acc Time
Random - 20.0 - 20.0 -
Just-LLM 7B 45.8 0.41 40.1 0.55
SF-VLM 13B 55.8 1.89 51.2 2.03
SOTA 20B 50.8 381 54.3 207
Therein, we introduce Multimodal Video Understanding (MVU), a simple framework that aggregates multimodal video-relevant information in an interpretable manner using natural language and achieves significant improvements over baselines across multiple datasets.